Iron Works International is considering a project that will produce annual cash flows of $38,500, $47,200, $57,900, and $23,400 over the next four years, respectively. What is the internal rate of return if the project has an initial cost of $112,500?
Answer #1

internal rate of return is the rate that makes initial investment equal to present value of cash inflows.

112,500 = 38,500 / (1 + r)1 + 47,200 / (1 + r)2 + 57,900 / (1 + r)3 + 23,400 / (1 + r)4

Using trial and error method, i.e., after trying various values for R, lets try R as 18.63%

112,500 = 38,500 / (1 + 0.1863)1 + 47,200 / (1 + 0.1863)2 + 57,900 / (1 + 0.1863)3 + 23,400 / (1 + 0.1863)4

112,500 = 32,453.8481 + 33,539.2184 + 34,681.2684 + 11,815.1091

112,500 = 112,500

Therefore, IRR is 18.63%